If the odds of winning a raffle are 7:229 what is the probability of winning?
a) 7/236
b) 229/236
c) 7/222
d) 7/229

Answer by ewatrrr(24785)


 
Hi,
odds of winning a raffle are 7:229
P(winning) = 7/(7+229) = 7/236